# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v1

@@ Abstract
Missing data often significantly hamper standard time series analysis, yet in practice they are frequently encountered. In this paper, we introduce temporal Wasserstein imputation, a novel method for imputing missing data in time series. Unlike existing techniques, our approach is fully nonparametric, circumventing the need for model specification prior to imputation, making it suitable for potential nonlinear dynamics. Its principled algorithmic implementation can seamlessly handle univariate or multivariate time series with any missing pattern. In addition, the plausible range and side information of the missing entries (such as box constraints) can easily be incorporated. As a key advantage, our method mitigates the distributional bias typical of many existing approaches, ensuring more reliable downstream statistical analysis using the imputed series. Leveraging the benign landscape of the optimization formulation, we establish the convergence of an alternating minimization algorithm to critical points. Furthermore, we provide conditions under which the marginal distributions of the underlying time series can be identified. Our numerical experiments, including extensive simulations covering linear and nonlinear time series models and an application to a real-world groundwater dataset laden with missing data, corroborate the practical usefulness of the proposed method.
Keywords— Missing data, Optimal transport, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
@@ 1 Introduction
In many scientific fields, data collection technology is often subject to various disruptions such as idiosyncratic or system-level instrument failures or changes in data-collection policies, leading to missing entries in the data (Little and Rubin,, 2019). In time series analysis, missing data pose serious challenges to the application of most statistical tools since it is typically presumed that observations are available at equally spaced time points. Moreover, since the dynamic relationship is of crucial importance in time series analysis, one can not proceed by throwing away the missing data, even if the missing data appear completely at random (i.e., the missing pattern is independent of the series).
To tackle this issue, one approach is to impute the missing entries, after which the vast array of time series methods can be readily applied for statistical inference in the downstream. Naturally, the estimation of the optimal interpolator, which is the expectation of the missing data conditional on the observed data, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021). Most of these methods are iterative or EM-type algorithms that alternate between model identification and missing data imputation (see, for example, Chpt. 4 of Peña and Tsay,, 2021). For certain time series models, one can estimate the model parameters simultaneously with the missing data (Harvey and Pierse,, 1984; Gómez et al.,, 1999). Besides the optimal interpolator, ad hoc deterministic imputation methods, such as smoothing with linear interpolation, B-splines and wavelets (see, for instance, Hastie et al.,, 2009), are also used in practice for their simplicity.
However, the optimal interpolator can introduce distributional biases that are detrimental to downstream statistical analysis. To illustrate, consider an AR(1) model, $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $10<|\phi|<$ 1 and $\{\epsilon_{t}\}$ is a sequence of i.i.d. standard normal random variables. It is well-known that the optimal interpolator can be expressed in terms of the inverse autocorrelation function (Peña and Maravall, 1991; see also Chpt. 6 of Peña et al.,, 2011). For instance, suppose that data at time $T$ and $T+1$ are missing; then the optimal interpolators are given by
$\hat{x}_{T}$ $=\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T+2},$
$\hat{x}_{T+1}$ $=\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T+2}.$
Using the fact that $\mathbb{E}(x_{t}x_{t-h})=\phi^{h}/(1-\phi^{2})$ for $h\geq 0$, it is not difficult to show
$\mathbb{E}(\hat{x}_{T}^{2})=\mathbb{E}(\hat{x}_{T+1}^{2})=\frac{\phi^{2}+3\phi^{4}+3\phi^{6}+2\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}\mathbb{E}(x_{t}^{2})<\mathbb{E}(x_{t}^{2}),$
and
@@
Hence, the variance and the autocovariances of the imputed data are different from those of the original AR(1) process. Intuitively, the optimal interpolator for consecutive missing data tends to predict the unconditional mean, leading to a downward bias (in magnitude) of the variances. In general, the joint distribution of the optimally interpolated time series is often quite different from that of the original process, leading to questionable downstream statistical analysis, especially when the number of missing data is non-negligible. Similar biases are also observed for some deterministic imputation methods (Carrizosa et al.,, 2013).
@@
In addition to the distributional bias, the above example also reveals further predicaments of the optimal interpolator. First, in order to compute the optimal interpolator, one often needs to choose a model class (such as the ARMA models) before imputation. However, such decision relies heavily on domain expertise and may be ill-informed when there are many missing values. Second, it may be difficult to compute the optimal interpolator if some nonlinear model is employed. With different missing patterns and different chosen models, the computational procedure may change drastically. Finally, incorporating side information about the missing data into this methodology appears less straightforward. Side information about the missing data restricts the set of admissible imputations. For instance, compositional time series is known to lie in $[0,1]$ and sums to unity at each time point. Besides the optimal interpolator, it is well-known that some deterministic interpolators such as B-spline smoothing are also inclined to over-(or under-)shoot, leading to imputations that are outside plausible ranges (Carrizosa et al.,, 2013).
@@
In this paper, we present a novel procedure for time series imputation, referred to as the temporal Wasserstein imputation (TWI), which circumvents the aforementioned limitations. As a nonparametric method, TWI is highly versatile and can be applied to a wide range of time series data. In particular, as a key advantage of TWI, it requires no model class specification prior to imputation, allowing it to perform well even when the underlying time series is nonlinear, such as the threshold autoregressive model (TAR, Tong,, 1983). Formulated as an optimization problem, TWI can be implemented efficiently through an alternating minimization algorithm. This principled algorithmic approach allows for the seamless integration of side information by restricting the admissible imputations, and its extension to multivariate time series is also straightforward. The optimization and statistical underpinnings of TWI are closely related to the optimal transport problem (Villani,, 2008, 2021; Cuturi,, 2013), which is briefly reviewed in Section 2. Section 3 maps out the optimization formulation and introduces the computational scheme in detail.
@@
The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the $p$-dimensional marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series (which is stable under stationarity), thereby lowering the distributional bias discussed above. In Section 4, we investigate the optimization and statistical properties of TWI. Our analysis shows that the optimization landscape of TWI is biconvex and the alternating algorithm is guaranteed to converge to a critical point. However, from the statistical viewpoint, missing data imputation suffers from an identification issue: Many distributions (of the stochastic process) may be equally plausible in describing the corrupted series. This raises the question of when and under what conditions TWI can identify the underlying distribution and yield imputations that can be safely used in downstream analysis. In Section 4, we show, via a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under assumptions on the missing pattern and the stability of the imputation. The latter condition motivates a simple modification of TWI, termed $k$-TWI, which shows further improvement in imputation quality from TWI in our numerical experiments.
@@
In Section 5, we evaluate the performance of TWI on synthetic data generated from a wide range of time series models, including nonlinear, multivariate, and nonstationary settings. Across all models, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics such as autocovariance and model parameter estimates. Its advantage is particularly prominent in the presence of nonlinear dynamics, where TWI consistently outperforms the benchmark methods. In Section 6, we apply TWI to a real-world groundwater dataset with many missing data. After imputation, we conduct factor analysis and autoregressive modeling in the downstream. Compared to existing methods, TWI effectively captures the underlying dynamics, producing more reasonable factor estimates as well as more accurate autoregressive predictions. These numerical results underscore the practical utility of TWI.
@@ 2 Preliminary: Optimal transport
In this section, we give a brief introduction to the optimal transport (OT) problem, which serves as a prerequisite to the rest of the paper. For a more comprehensive account on the theory of optimal transport, the readers are referred to Villani, (2021, 2008).
Given two probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{p}$, and a cost function $L:\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}_{+}$, the optimal transport cost between $\mu$ and $\nu$ is defined as
@@
If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$ for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we formally denote as $\mathcal{W}_{k}(\mu,\nu)$. Namely,
@@
The Wasserstein distance is a metric on the space of probabilities with finite $k$-th moment on $\mathbb{R}^{p}$, and metrizes the weak topology on this space (Villani,, 2021, Chpt. 9). In particular, if $\mu$ and $\{\mu_{t}:t=1,2,\ldots\}$ are probability measures on $\mathbb{R}^{p}$ such that $\mathcal{W}_{k}(\mu_{t},\mu)\rightarrow 0$ as $t\rightarrow\infty$, then $\mu_{t}\Rightarrow\mu$, where $\Rightarrow$ denotes the weak convergence of probability measures (Billingsley,, 1995). Therefore, it can be used as a discrepancy metric between $\mu$ and $\nu$. In fact, the Wasserstein distance offers several advantages over other discrepancy metrics such as the Kullbeck-Leibler divergence or the Kolmogorov-Smirnov distance. In addition to its metric properties, the Wasserstein distance has a nice dual representation and handles support mismatch easily, which makes it suitable in modern machine learning and statistical applications.
@@
$\mathsf{OTC}(\mu,\nu)=\min_{\{\pi_{ij}\}}$ $\sum_{i,j}L(\mathbf{x}_{i},\mathbf{y}_{j})\pi_{ij}$
s.t. $\sum_{i=1}^{n}\pi_{ij}=q_{j},\quad j=1,\ldots,m,$ (1)
$\sum_{j=1}^{m}\pi_{ij}=p_{i},\quad i=1,\ldots,n,$ (2)
$\pi_{ij}\geq 0.$
The constraints (1) and (2) enforce the joint distribution $\{\pi_{ij}\}$ to be a coupling between $\mu$ and $\nu$, and will henceforth be called the coupling constraints. In the discrete case, the optimal coupling $\{\pi_{ij}\}$, in general, cannot be expressed as the push-forward under an one-to-one map. Instead, the optimal coupling typically involves split of masses for discrete measures. Many computational methods have been proposed to solve this linear program, including the network flow simplex (Orlin,, 1997) and the auction method (Bertsekas,, 1992). When the problem dimensions $n,m$ are large, standard tools for linear programs may suffer from slow computation. Nevertheless, recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale parallelization and can be accelerated with GPUs; see Cuturi, (2013) and Peyré and Cuturi, (2019).
@@ 3 Temporal Wasserstein imputation
In this section, we introduce temporal Wasserstein imputation (TWI), the proposed method for time series imputation, through an optimal transport formulation. First, an optimization formulation of the proposed method that directly connects to the optimal transport problem is laid out. Then, we describe an alternating minimization algorithm to impute the missing values. The algorithm also reveals some connections between the proposed method and existing approaches.
@@ 3.1 Optimization formulation
Let $\{\tilde{x}_{t}\}$ be the time series of interest, and we observe the time series $\{x_{t}:t=0,1,\ldots,n-1\}$. Denote by $\bigstar$ a generic missing value, and $\mathcal{M}$ the set of time indices corresponding to the missing observations. Then, $x_{t}=\bigstar$ if $t\in\mathcal{M}$ and $x_{t}=\tilde{x}_{t}$ otherwise. If $\{\tilde{x}_{t}\}$ is stationary, it follows that its $p$-dimensional marginal distributions, i.e. the probability distributions of $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$, are stable over time. Let $\mu_{p}$ ($p\in\mathbb{N}$) be the probability measure associated with $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$. Stationarity implies $\mu_{p}$ is independent of the time index $t$.
In light of this property, when imputing the missing values in $(x_{0},\ldots,x_{n-1})$, we seek an imputation $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n-1})^{\top}$ whose marginal distributions remain nearly unchanged before and after a specified time point $n_{1}$. More specifically, we first divide the time series into two parts $\{w_{0},w_{1},\ldots,w_{n_{1}}\}$ and $\{w_{n_{1}+1},w_{n_{1}+2},\ldots,w_{n-1}\}$, where $p-1<n_{1}<n-p$, and define the two empirical marginal distributions
$\hat{\mu}_{\mathrm{pre}}(\mathbf{w})=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})},$
$\hat{\mu}_{\mathrm{post}}(\mathbf{w})=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}.$
The discrepancy between the two empirical distributions can then be measured by the optimal transport cost, which is defined as the minimum value of the following linear program.
$\begin{split}\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij}}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t. }&\sum_{i=p-1}^{n_{1}}\pi_{ij}=\frac{1}{n-n_{1}-1}\\ &\sum_{j=n_{1}+1}^{n-1}\pi_{ij}=\frac{1}{n_{1}-p+2}\\ &\pi_{ij}\geq 0\mbox{ for all $i,j$},\end{split}$ (3)
where $\mathbf{v}_{i}(\mathbf{w}):=(w_{i},w_{i-1},\ldots,w_{i-p+1})^{\top}$, and, for ease of notation, we enumerate the rows and columns of the $(n_{1}-p+2)\times(n-n_{1}-1)$ matrix $\mathbf{\Pi}$ by $\{p-1,p,\ldots,n_{1}\}$ and $\{n_{1}+1,\ldots,n-1\}$, respectively. An important special case of (3) is when $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{k}$ for some $k\geq 1$. In this case, we have $\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}},\hat{\mu}_{\mathrm{post}})=\mathcal{W}_{k}^{k}(\hat{\mu}_{\mathrm{pre}},\hat{\mu}_{\mathrm{post}})$.
@@
where the constraint set $\mathcal{C}$ delineates the admissible imputations. Typically, it enforces the imputation to agree with the observed data, in which case $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$. It may also reflect the side information available. For example, if $x_{t}$ are weekly sales data, the knowledge of monthly sales define a set of linear constraints for the missing values. Another example is when $\{\mathbf{x}_{t}\}$ is a $d$-dimensional compositional time series such as the market share of several companies in an industry. It follows that $\mathbf{x}_{t}^{\top}\mathbf{1}=1$ for all $t$. Then, by defining $\mathcal{C}=\{\mathbf{W}\in\mathbb{R}^{n\times d}:\mathbf{W}\mathbf{1}_{d}=\mathbf{1}_{n}\}$, where $\mathbf{1}_{n}=(1,\ldots,1)^{\top}\in\mathbb{R}^{n}$, the set of admissible imputations are restricted to be compositional as well.
@@
where $\lambda>0$ is small. As we will see in the next subsection, the $\ell_{2}$-regularization term makes a subproblem in the alternating minimization algorithm strongly convex, enhancing numerical stability. In addition, in the important special case corresponding to the Wasserstein distance of order 2, i.e., $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{2}$, the regularization allows one to solve the subproblem by finding a fixed point of a linear system. Since $L$ is often chosen to be some power of the Euclidean norm, which induces the Wasserstein distances between the marginal distributions of a time series, the solution $\hat{\mathbf{w}}$ to (4) or (5) is called the temporal Wasserstein imputation (TWI) throughout this paper.
In this formulation, it is clear that the temporal Wasserstein imputation, which seeks to equate the marginal distributions on both sides of $n_{1}$, is nonparametric. In addition, with a mildly large $p$, the $p$-dimensional marginal distributions can fully capture the nonlinear dynamics of a Markov process (such as a finite-order TAR model), which would otherwise need to be approximated using many parameters in an ARMA model.

### v2

@@ Abstract
Missing data can significantly hamper standard time series analysis, yet in practice they are frequently encountered. In this paper, we introduce temporal Wasserstein imputation, a novel method for imputing missing data in time series. Unlike existing techniques, our approach is fully nonparametric, circumventing the need for model specification prior to imputation, making it suitable for potential nonlinear dynamics. Its principled algorithmic implementation can seamlessly handle univariate or multivariate time series with any non-systematic missing pattern. In addition, the plausible range and side information of the missing entries (such as box constraints) can easily be incorporated. As a key advantage, our method mitigates the distributional bias typical of many existing approaches, ensuring more reliable downstream statistical analysis using the imputed series. Leveraging the benign landscape of the optimization formulation, we establish the convergence of an alternating minimization algorithm to critical points. We also provide conditions under which the marginal distributions of the underlying time series can be identified. Numerical experiments, including extensive simulations covering linear and nonlinear time series models and a real-world groundwater dataset laden with missing values, corroborate the practical usefulness of the proposed method.
Keywords: Optimal transport, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
@@ 1 Introduction
In many scientific fields, data collection is often subject to various disruptions such as idiosyncratic or system-level instrument failures or changes in data collection policies, leading to missing entries in the data (Little and Rubin,, 2019). In time series analysis, missing data pose serious challenges to the application of many statistical tools since it is typically presumed that observations are available at equally spaced time points. Moreover, since the dynamic relationship is of crucial importance in time series analysis, one cannot proceed by throwing away the missing data, even if the missing data appear completely at random (i.e., the missing pattern is independent of the series).
To tackle this issue, one approach is to impute the missing entries, after which the vast array of time series methods becomes readily applicable for statistical inference in the downstream. Naturally, the estimation of the optimal interpolator, which is the expectation of the missing data conditional on the observed data, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021; McElroy and Politis,, 2020, 2022; McElroy,, 2022). Since the optimal interpolator would depend on the probabilistic structure of the underlying process, most methods involve EM-type iterations that alternate between model identification and missing data imputation (see, for example, Chpt. 4 of Peña and Tsay,, 2021). Besides the optimal interpolator, ad hoc deterministic imputation methods, such as smoothing with linear interpolation, B-splines and wavelets (see, for instance, Hastie et al.,, 2009), are also used in practice for their simplicity.
However, the optimal interpolator can cause distributional biases that are detrimental to downstream statistical analysis. To illustrate, consider an AR(1) process, $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $10<|\phi|<$ 1 and $\{\epsilon_{t}\}$ is a sequence of i.i.d. standard normal random variables. The optimal interpolator can be expressed in terms of the inverse autocorrelation functions (Peña and Maravall, 1991; McElroy and Politis, 2022; see also Chpt. 6 of Peña et al.,, 2011). For instance, suppose that data at time indices $T$ and $T+1$ are missing. Then the optimal interpolators are given by
$\hat{x}_{T}=\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T+2},\quad\hat{x}_{T+1}=\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T+2}.$
Using the fact that $\mathbb{E}(x_{t}x_{t-h})=\phi^{h}/(1-\phi^{2})$, for $h\geq 0$, it is not difficult to show that
$\mathbb{E}(\hat{x}_{T}^{2})=\mathbb{E}(\hat{x}_{T+1}^{2})=\frac{\phi^{2}+3\phi^{4}+3\phi^{6}+2\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}\mathbb{E}(x_{t}^{2})<\mathbb{E}(x_{t}^{2}),\mbox{ and}$
@@
Hence, the variance and the autocovariances of the imputed data are different from those of the original AR(1) process. In general, the joint distribution of the optimally interpolated time series is often quite different from that of the original process, leading to questionable downstream statistical analysis, especially when the number of missing data is non-negligible. Similar biases are also observed for some deterministic imputation methods (Carrizosa et al.,, 2013).
@@
In addition to the distributional bias, the above example also reveals further predicaments of the optimal interpolator. First, in order to compute the optimal interpolator, one often needs to choose a model class before imputation. However, such a decision relies heavily on domain expertise and may be ill-informed when there are many missing values. While it is possible, for certain time series models, to estimate the model parameters without imputation, it still engages in model specification before imputation. Second, it may be difficult to compute the optimal interpolator if some nonlinear model is employed. With different missing patterns and different chosen models, the computational procedure may change drastically. Finally, incorporating side information about the missing data into this methodology appears less straightforward. Side information about the missing data restricts the set of admissible imputations. For instance, a compositional time series is known to lie in $[0,1]$ and sums to unity at each time point. Besides the optimal interpolator, many deterministic interpolators such as B-spline smoothing are also inclined to over-(or under-)shoot, leading to imputations that are outside plausible ranges.
@@
In this paper, we present the temporal Wasserstein imputation (TWI), a novel approach for time series imputation which circumvents the aforementioned limitations. As a nonparametric method, TWI is highly versatile and can be applied to a wide range of time series data. In particular, as a key advantage of TWI, it requires no model class specification prior to imputation, allowing it to perform well even when the underlying time series is nonlinear, such as the threshold autoregressive processes (TAR, Tong,, 1983). Formulated as an optimization problem, TWI can be implemented efficiently through an alternating minimization algorithm. This principled algorithmic approach allows for the seamless integration of side information by restricting the admissible imputations, and its extension to multivariate time series is also straightforward. The optimization and statistical underpinnings of TWI are closely related to the optimal transport problem (Villani,, 2008, 2021; Cuturi,, 2013), which is briefly reviewed in Section 2. Section 3 maps out the optimization formulation and introduces the computational scheme in detail.
@@
The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series, thereby lowering the distributional bias discussed above. In Section 4, we investigate the optimization and statistical properties of TWI. Our analysis shows that the optimization landscape of TWI is biconvex and the alternating algorithm is guaranteed to converge to a critical point. However, from the statistical viewpoint, missing data imputation suffers from an identification issue: Many distributions (of the stochastic process) may be equally plausible in describing the corrupted series. This raises the question of whether and under what conditions TWI can identify the underlying marginal distributions. In Section 4.2, we show, via a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under assumptions on the missing pattern and the stability of the imputation. The latter condition motivates a simple modification of TWI, termed $k$-TWI, which shows further improvement in imputation quality from TWI in our numerical experiments.
@@
In Section 5, we evaluate the performance of TWI on synthetic data generated from a wide range of time series models, including nonlinear, multivariate, and nonstationary settings. Across all processes considered, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics such as autocovariance and model parameter estimates. Its advantage is particularly prominent in the presence of nonlinear dynamics. In Section 6, we apply TWI to a real-world groundwater dataset with many missing data. After imputation, we conduct independent component analysis and autoregressive modeling in the downstream. Compared to existing methods, TWI effectively captures the underlying dynamics, producing interpretable components as well as accurate autoregressive predictions. These numerical results support the practical contributions of TWI.
@@ 2 Preliminary: Optimal transport
In this section, we briefly review the optimal transport (OT) problem, which serves as a prerequisite to the rest of the paper. For a more comprehensive account on the theory of optimal transport, the readers are referred to Villani, (2021, 2008).
Given two probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{p}$ and a cost function $L:\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}_{+}$, the optimal transport cost between $\mu$ and $\nu$ is defined as
@@
If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$, for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we formally denote as $\mathcal{W}_{k}(\mu,\nu)$. Namely,
@@
The Wasserstein distance is a metric on the space of probabilities with finite $k$-th moment on $\mathbb{R}^{p}$, and metrizes the weak topology on this space (Villani,, 2021, Chpt. 9). In particular, if $\mu$ and $\{\mu_{t}:t=1,2,\ldots\}$ are probability measures on $\mathbb{R}^{p}$ such that $\mathcal{W}_{k}(\mu_{t},\mu)\rightarrow 0$ as $t\rightarrow\infty$, then $\mu_{t}\Rightarrow\mu$ and $\int\|\mathbf{x}\|^{k}\mu_{t}(d\mathbf{x})\rightarrow\int\|\mathbf{x}\|^{k}\mu(d\mathbf{x})$, where $\Rightarrow$ denotes the weak convergence of probability measures (Billingsley,, 1995). Therefore, it can be used as a discrepancy metric between $\mu$ and $\nu$. In fact, the Wasserstein distance offers several advantages over other discrepancy metrics such as the Kullbeck-Leibler divergence and the Kolmogorov-Smirnov distance. For instance, the Wasserstein distance has a nice dual representation and handles support mismatch easily, making it suitable for machine learning and statistical applications.
@@
$\mathsf{OTC}(\mu,\nu)=\min_{\{\pi_{ij}\},\pi_{ij}\geq 0}$ $\sum_{i,j}L(\mathbf{x}_{i},\mathbf{y}_{j})\pi_{ij}$
s.t. $\sum_{i=1}^{n}\pi_{ij}=q_{j},\quad j=1,\ldots,m,$ $\sum_{j=1}^{m}\pi_{ij}=p_{i},\quad i=1,\ldots,n.$ (1)
The constraints in (1) enforce the joint distribution $\{\pi_{ij}\}$ to be a coupling between $\mu$ and $\nu$, and will henceforth be called the coupling constraints. In the discrete case, the optimal coupling $\{\pi_{ij}\}$ typically involves split of masses for discrete measures and is no longer the push-forward under an one-to-one map. Recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale parallelization and can be accelerated with GPUs; see Cuturi, (2013) and Peyré and Cuturi, (2019).
@@ 3 Temporal Wasserstein imputation
In this section, we introduce the proposed temporal Wasserstein imputation (TWI) method for time series imputation. First, we provide an optimization formulation of the proposed method that directly connects to the optimal transport problem. Then, we describe an alternating minimization algorithm to impute the missing values. The algorithm also reveals some connections between the proposed method and some existing approaches.
@@ 3.1 Optimization formulation
Let $\{\tilde{x}_{t}\}$ be the time series of interest and we observe the time series $\{x_{t}:t=0,1,\ldots,n-1\}$. Denote by $\bigstar$ a generic missing value, and $\mathcal{M}$ the set of time indices corresponding to the missing observations. Thus, $x_{t}=\bigstar$ if $t\in\mathcal{M}$ and $x_{t}=\tilde{x}_{t}$ otherwise. We discuss below the case that $x_{t}$ is stationary. The case of unit-root nonstationary series is discussed at the end of this section. If $\{\tilde{x}_{t}\}$ is stationary, then its $p$-dimensional marginal distribution $\mu_{p}$, i.e. the distributions of $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$, does not depend on $t$, where $p\in\mathbb{N}$. In view of this property, when imputing the missing values in $(x_{0},\ldots,x_{n-1})$, we seek an imputation $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n-1})^{\top}$ whose marginal distribution remains nearly unchanged. More specifically, we first split the time series into two parts $\{w_{0},w_{1},\ldots,w_{n_{1}}\}$ and $\{w_{n_{1}+1},w_{n_{1}+2},\ldots,w_{n-1}\}$ and define the two empirical marginal distributions
$\hat{\mu}_{\mathrm{pre}}(\mathbf{w})=\sum_{t=p-1}^{n_{1}}\frac{\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}}{n_{1}-p+2},\quad\hat{\mu}_{\mathrm{post}}(\mathbf{w})=\sum_{t=n_{1}+1}^{n-1}\frac{\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}}{n-n_{1}-1},$
where $p-1<n_{1}<n-p$, and the values of $p$ and $n_{1}$ used will be discussed later. The discrepancy between the two empirical distributions can then be measured by the optimal transport cost, which is the minimum value of the following linear program.
$\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij},\pi_{ij}\geq 0}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t. }&\sum_{i=p-1}^{n_{1}}\pi_{ij}=\frac{1}{n-n_{1}-1}\quad\mbox{for all }j\\ &\sum_{j=n_{1}+1}^{n-1}\pi_{ij}=\frac{1}{n_{1}-p+2}\quad\mbox{for all }i,\end{aligned}\right\}$ (2)
where $\mathbf{v}_{i}(\mathbf{w}):=(w_{i},w_{i-1},\ldots,w_{i-p+1})^{\top}$, and, for ease of notation, we enumerate the rows and columns of the $(n_{1}-p+2)\times(n-n_{1}-1)$ matrix $\mathbf{\Pi}$ by $\{p-1,p,\ldots,n_{1}\}$ and $\{n_{1}+1,\ldots,n-1\}$, respectively. An important special case of (2) is when $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{k}$, for some $k\geq 1$. In this case, we have $\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}},\hat{\mu}_{\mathrm{post}})=\mathcal{W}_{k}^{k}(\hat{\mu}_{\mathrm{pre}},\hat{\mu}_{\mathrm{post}})$.
@@
where the constraint set $\mathcal{C}$ delineates the admissible imputations. Typically, it enforces the imputation to agree with the observed data, in which case $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$. It may also reflect the side information available. For example, if $x_{t}$ are monthly sales data, the knowledge of quarterly sales define a set of linear constraints for the missing values. Another example is when $\{\mathbf{x}_{t}\}$ is a $d$-dimensional compositional time series such as the market share of several companies in an industry. It follows that $\mathbf{x}_{t}^{\top}\mathbf{1}=1$ for all $t$. Then, by defining $\mathcal{C}=\{\mathbf{W}\in\mathbb{R}^{n\times d}:\mathbf{W}\mathbf{1}_{d}=\mathbf{1}_{n}\}$, where $\mathbf{1}_{n}=(1,\ldots,1)^{\top}\in\mathbb{R}^{n}$, the set of admissible imputations are restricted to be compositional as well.
@@
where the regularization parameter $\lambda>0$ is small. As we will see in the next subsection, the $\ell_{2}$-regularization term makes a subproblem in the alternating minimization algorithm strongly convex, enhancing numerical stability. In addition, in the important special case corresponding to the Wasserstein distance of order 2, i.e., $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{2}$, the regularization allows one to solve the subproblem by finding a fixed point of a linear system. Since $L$ is often chosen to be some power of the Euclidean norm, which induces the Wasserstein distances between the marginal distributions of a time series, the solution $\hat{\mathbf{w}}$ to (3) or (4) is called the temporal Wasserstein imputation (TWI) throughout this paper.
In this formulation, it is clear that the temporal Wasserstein imputation, which seeks to equate the marginal distributions on both sides of $n_{1}$, is nonparametric. With a mildly large $p$, the $p$-dimensional marginal distributions can sometimes fully characterize the nonlinear dynamics of a Markov process (such as a finite-order threshold AR process), which would otherwise need many parameters using ARMA approximation.